OpenStudy (anonymous):

Find the coordinates of the point on the curve y = x^2 -x -1 where the tangent is perpendicular to the tangent at the point (-1,1)

OpenStudy (anonymous):

@redpellet \[y'(x) = \frac 1 3\\ x =\frac 23\\ \left(\frac 23,-\frac {11}9\right) \] is the point

OpenStudy (anonymous):

For the slope of the tangent at the point (-1, 1): y = x^2 - x - 1 y' = 2x - 1 y' = 2(-1) - 1 = -3 therefore its slope is -3. (a perpendicular slope is the negative reciprocal): so the slope of line perpendicular to the tangent is 1/3. 1/3 = 2x - 1 x = 2/3 and then sub that into y = x^2 -x -1 for the y component of the coordinate the point on the graph at that x value is where the slope of its tangent is perpendicular to the tangent at (-1, 1)